OpenStudy (anonymous):

The isotope Cobalt-60 has a nuclear mass of 59.933820 u 3. Calculate the Binding Energy per Nucleon of Cobalt-60 using the following information. Mass of Proton: 1.007825 u Mass of Neutron: 1.008665 u 1 u = 931.5 MeV Please help! If you can show the answer and the work, with an explanation afterwards, that would be great although you don't have to I just want to be able to understand it - thanks!

OpenStudy (anonymous):

Okay so the way to solve "Binding Energy Per Nucleon" is similar to regular binding energy! Find your defect mass by doing Total Mass=(27p*1.00725)+(33n*1.008665) Total Mass=(27.211275)+(33.285945) Total Mass=60.49722 Defect Mass=60.49722-59.93382 Defect Mass=.5644 Then almost the same way to find binding energy, Binding Energy Per Nucleon=.5644*931.5/60 Binding Energy Per Nucleon=525.7386/60 Binding Energy Per Nucleon=8.76231 MeV Hope this helped!